A firm has the following revenue and cost functions.

TR = 120 Q – Q2

TC = 12 Q2 +30 Q + 10

Determine the quantity level at which the firm maximizes its total profit. 



1
Expert's answer
2021-07-20T10:42:36-0400

Marginal revenue:

MR  "=\frac {d}{dQ}" "(TR)" "=" "\frac{d}{dQ}" "(" 120Q"-" Q"^2" ")"

"=" 120"-" 2Q

Marginal Cost:

"MC" "=" "\frac{d}{dQ}" ("TC" ")" "=" "\frac{d}{dQ}" "(" 12Q"^2" "+" 30Q"+" 30")" "=" 24Q"+" 30


Firm maximizes total profit when:

MR "=" MC

120"-" 2Q"=" 24Q"+" 30

26Q"=" 90

Q"=" 3.46